OpenStudy (anonymous):

how do u simplify (x-3)/(4x^2) - (2x+2)/(9x) + (7x)/24

OpenStudy (amistre64):

aparently by getting someone else to do it for you lol ;)

OpenStudy (anonymous):

i know

OpenStudy (amistre64):

get like denominators and drudge thru it.... its dirty work but build character :)

OpenStudy (anonymous):

i need some steps to do the work. do know how to do it

OpenStudy (amistre64):

(x-3)/(4x^2) - (2x+2)/(9x) + (7x)/24...i do

OpenStudy (anonymous):

can u hlp me

OpenStudy (amistre64):

x-3 2x+2 7x ---- - ----- + --- lets start buy re writing the numbers... 4x^2 9x 24

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

we need to simply mulitply everything in the bootom to get a common denom right? 4x^2( 9x)( 24) = 864x^3 correct?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

now just multiply everything by that number really; or rather it might be easier to see what happens when we leave it in its factored form of: (4x^2)(9x)(24) (x-3) * (4x^2)(9x)(24) ------------------- = ? (4x^2) (x-3)(9x)(24) right? which equals 216x^2 - 648x

OpenStudy (amistre64):

(2x+2) * (4x^2)(9x)(24) --------------------- (9x) which equals 192x3 +192x^2 right?

OpenStudy (amistre64):

7x (4x^2)(9x)(24) --------------- (24) 7x (4x^2)(9x) = 252x^3

OpenStudy (amistre64):

now lets put it all together 192x^3 +192x^2 252x^3 +216x^2 - 648x ----------------------- 444x^3 + 408x^2 - 648x; maybe wanna factor that ...and see if i did any errors

OpenStudy (anonymous):

no their r no errors

OpenStudy (amistre64):

444x^3 408x^2 648x ------- + ------- - ------- is gonna be our answer i think 864x^3 864x^3 864x^3

OpenStudy (amistre64):

i think we can do that with an easier number tho; we dint use the "least" common denomintor.....just a common one

OpenStudy (anonymous):

I'm starding to understand it now. I have another question do u know how to find the conic section of 3y^2+9x-8y=13

OpenStudy (amistre64):

yes; its a parabola

OpenStudy (amistre64):

3y^2 -8y = 13- 9x complete the square on the y side...

OpenStudy (anonymous):

how do u know

OpenStudy (amistre64):

its only got one ^2 part to it; so it bends only once

OpenStudy (amistre64):

the rest of it is just filler to make it bend how and where you want it

OpenStudy (anonymous):

Oh, thanks 4 the help

OpenStudy (amistre64):

youre welcome :)